Identify the graph of the conic section defined by the following equation.
x^2+4x-y+9=0

Answer by lwsshak3(11628)

Identify the graph of the conic section defined by the following equation.
x^2+4x-y+9=0
**
x^2+4x-y+9=0
y=x^2+4x+9
complete the square
y=(x^2+4x+4)+9-4
y=(x+2)^2+5
This is an equation of a parabola of the standard form: y=A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex.
For given equation:
Vertex: (-2,5)
Parabola opens upwards, has a minimum
Axis of symmetry: x=-2